# Curvature-Constrained Vector Field for Motion Planning of Nonholonomic Robots

### v1

@@ VI-B3 Results
To evaluate the convergence towards the target positive limit set and the tracking performance of 3D setpoints, we define the distance error between the current position and the target positive limit set as $d_{e}=r_{\delta}-r_{2}$ and present the attitude error with $\|\log_{\mathrm{SO}(3)}{\bm{R}_{e}}\|$ where $\bm{R}_{e}=\bm{R}\bm{R}_{r}^{-1}$ and $\bm{R}_{r}=\bm{R}_{z}(\alpha_{d})\bm{R}_{y}(\beta_{d})\bm{R}_{z}(\gamma_{d})$. The results for the HIL experiments on fixed-wing UAV are presented in Figures 10 and 11.
Despite the error introduced by the simplification in the derivation of setpoints, dynamics uncertainties and measurement noises, the control inputs can be effectively tracked by the fixed-wing UAV. This is because the proposed algorithm addresses the curvature constraint and the nonholonomic constraint, which are the most critical kinematic constraints for the UAV. Consequently, both the distance error $d_{e}$ and the attitude error $\|\log_{\mathrm{SO}(3)}{\bm{R}_{e}}\|$ in Figures 11 are stabilized as soon as the motion planning algorithm is activated. This confirms the convergence of the UAV to the target positive limit set and the arrival at the target configuration under curvature constraint as shown in Figure 10. Moreover, the results also demonstrate that the proposed algorithm can be implemented onboard in real-time, and can handle time-varying tasks where the target configuration switches during the flight, which is another merit of the closed-form motion planning method.
@@ Appendix A Proof of Lemma 1
as provided in [41, Chapter 13.3]. We utilize the cross product in $\mathbb{R}^{2}$ by defining $[a,b]\times[c,d]\coloneq ad-bc$.
When the VF is given in the polar coordinates, the integral curve is the solution to the following ordinary differential equation:
$\left[\begin{matrix}\dot{r}\\ r\dot{\varphi}\end{matrix}\right]=\left[\begin{matrix}F_{r}\\ F_{\varphi}\end{matrix}\right].$ (32)
Therefore, each component in (31) can be computed as follows:
$\bm{p}$ $=\left[\begin{matrix}r\cos{\varphi}\\ r\sin{\varphi}\end{matrix}\right]=\exp(\varphi^{\wedge})\left[\begin{matrix}r\\ 0\end{matrix}\right],$ (33)
$\dot{\bm{p}}$ $=\exp(\varphi^{\wedge})\left[\begin{matrix}F_{r}\\ F_{\varphi}\end{matrix}\right],$
$\ddot{\bm{p}}$ $=\frac{\mathrm{d}}{\mathrm{d}t}(\exp(\varphi^{\wedge}))\left[\begin{matrix}F_{r}\\ F_{\varphi}\end{matrix}\right]+\exp(\varphi^{\wedge})\left[\begin{matrix}\dot{F_{r}}\\ \dot{F_{\varphi}}\end{matrix}\right]$
$=\exp(\varphi^{\wedge})\left[\begin{matrix}\frac{\partial{F_{r}}}{\partial{r}}&\frac{\partial{F_{r}}}{r\partial{\varphi}}-\frac{F_{\varphi}}{r}\\ \frac{\partial{F_{\varphi}}}{\partial{r}}+\frac{F_{\varphi}}{r}&\frac{\partial{F_{\varphi}}}{r\partial{\varphi}}\end{matrix}\right]\mathbf{F}.$
By substituting (33) into (31), the result is equivalent to (4) and the proof is completed.
@@ Appendix B Proof of Theorem 1
[…]
@@
Given (13), it can be verified that $\mathbf{F}(\bm{p})\in\mathcal{C}_{\mathcal{A}_{inv}}(\bm{p})$, $\forall\bm{p}\in\partial\mathcal{A}_{inv}$, implying $\mathcal{A}_{inv}$ is positively invariant according to Nagumo’s theorem [42]. Additionally, the positively invariant annular region $\mathcal{A}_{inv}$ contains at least one limit cycle since it contains no singular point and $\mathbf{F}(\bm{p}),\bm{p}\right>\neq 0$, $\forall\bm{p}\in\partial\mathcal{A}_{inv}$ indicates the existence of forward orbit from the boundary [39, Theorem 1.179].
@@
According to (34a), there hold $\dot{r}_{d}>0$ for $-r_{2}<r_{d}<0$ and $\dot{r}_{d}<0$ for $r_{d}>0$. Subsequently, we choose the positive definite Lyapunov function candidate
@@
3) Since the VF $\mathbf{F}$ is of class $C^{1}$, we can conclude that the curvature is continuous over $\mathcal{D}$ according to Lemma 1. To show the boundedness of the curvature, we firstly discuss the trivial case where $\bm{p}\in\mathcal{A}_{1}\setminus\bm{0}$. In this case, there hold $F_{r}=1$ and $F_{\varphi}=0$ and $\bm{K}_{F}=\bm{0}$. Substituting these values into (4), we have $\kappa_{\mathbf{F}}(\bm{p})=0$, $\forall\bm{p}\in\mathcal{A}_{1}\setminus\bm{0}$. Similarly we can obtain the same result that $\kappa_{\mathbf{F}}(\bm{p})=0$, $\forall\bm{p}\in\mathcal{A}_{4}$, where $F_{r}=-1$ and $F_{\varphi}=0$.
Next, consider $\bm{p}\in\mathcal{A}_{2}$, where $F_{r}=\lambda(s_{2})$, $F_{\varphi}=\tilde{\lambda}(s_{2})$, $\lambda(s_{2})=2s_{2}^{3}-3s_{2}^{2}+1$ and $s_{2}=(r-r_{1})(r_{2}-r_{1})\in[0,1)$. Substituting the equations above into (4) yields
@@
[…]
@@
[…]
@@ Appendix C Proof of Theorem 2
In this proof, we show that the proposed CVF has a bounded and continuous curvature, the target position $\bm{p}_{d}$ belongs to the positive limit set of integral curves and $\mathbf{T}(\bm{p}_{d})/\|\mathbf{T}(\bm{p}_{d})\|=[\cos{\theta_{d}},\sin{\theta_{d}}]^{\intercal}$.
1) Upper bounded continuous curvature
According to [39, Chapter 1.5], the normalization operation reparameterizes the time of the integral curves of $\mathbf{T}$ from $\mathbf{F}$. In other words, the shape of integral curves of $\mathbf{T}$ with $\bm{\zeta}_{0}=\bm{p}$ remains the same as those of $\mathbf{F}$ with $\bm{\zeta}_{0}=\bm{p}+\bm{p}_{\delta}$. Therefore, the integral curves from $\mathbf{F}$ to $\mathbf{T}$ are translated by $\bm{p}_{\delta}$. In addition, the translation is a positive isometry, which preserves the curvature of the integral curves [43]. Therefore, the integral curves of $\mathbf{T}$ still have continuous curvature bounded by $\bar{\kappa}$ as shown in Theorem 1.
2) Positive limit set of integral curves of $\mathbf{T}$
@@
3) $\mathbf{T}(\bm{p}_{d})$ aligning with $\theta_{d}$
Given (15), we transform the components of $\mathbf{F}$ from $(r,\varphi)$ coordinates into Cartesian coordinates by rotation matrix $\exp(\varphi^{\wedge})$ and there holds that
@@
$=\lim_{\begin{subarray}{c}r\to r_{2}\\ \varphi\to\theta_{d}-\frac{\pi}{2}\end{subarray}}\exp(\varphi^{\wedge})\left[\begin{matrix}F_{r}(r)\\ F_{\varphi}(r)\end{matrix}\right]/\|\exp(\varphi\wedge)\left[\begin{matrix}F_{r}(r)\\ F_{\varphi}(r)\end{matrix}\right]\|$
@@ Appendix D Kinematics of Reference Orientation
[…]
@@
$\left[\begin{matrix}T_{x}\\ T_{y}\end{matrix}\right]=\exp{(\varphi_{\delta}^{\wedge})}\left[\begin{matrix}T_{r_{\delta}}\\ T_{\varphi_{\delta}}\end{matrix}\right],$
@@
$\left[\begin{matrix}-T_{y}&T_{x}\end{matrix}\right]=\left[\begin{matrix}T_{r_{\delta}}&T_{\varphi_{\delta}}\end{matrix}\right]\exp{\left(-(\varphi_{\delta}+\frac{\pi}{2})^{\wedge}\right)}.$
@@
$\dot{\mathbf{T}}=\left[\begin{matrix}\dot{T}_{x}\\ \dot{T}_{y}\end{matrix}\right]=\frac{\mathrm{d}}{\mathrm{d}t}\left(\exp{(\varphi_{\delta}^{\wedge})}\left[\begin{matrix}T_{r_{\delta}}\\ T_{\varphi_{\delta}}\end{matrix}\right]\right)$
$=\frac{\mathrm{d}\exp{(\varphi_{\delta}^{\wedge})}}{\mathrm{d}t}\left[\begin{matrix}T_{r_{\delta}}\\ T_{\varphi_{\delta}}\end{matrix}\right]+\exp{(\varphi_{\delta}^{\wedge})}\frac{\mathrm{d}}{\mathrm{d}t}\left[\begin{matrix}T_{r_{\delta}}\\ T_{\varphi_{\delta}}\end{matrix}\right].$
@@
where $A(r_{\delta})=\sqrt{1/r_{\delta}^{2}+\theta_{r,r_{\delta}}^{2}}$ is the magnitude of the gradient $\nabla\theta_{r}$, $\Delta\theta=\theta-\theta_{grad}$ and $\theta_{grad}=\varphi_{\delta}+\mathrm{arctan}(1/(r_{\delta}\theta_{r,r_{\delta}}))$ specifies the orientation of $\nabla\theta_{r}$. The partial derivative of $\theta_{r}$ w.r.t. $r_{\delta}$ is given by
@@
where $s_{i}=(r-r_{i-1})/(r_{i}-r_{i-1})$, $\lambda_{i}=2s^{3}_{i}-3s^{2}_{i}+1$ for $i=2,3$ and $\mathcal{A}_{i}$ are defined in (9).

### v2

@@ VI-B3 Results
To evaluate the convergence towards the target positive limit set and the tracking performance of 3D setpoints, we define the distance error between the current position and the target positive limit set as $d_{e}=r_{\delta}-r_{2}$ and present the attitude error with $\|\log_{\mathrm{SO}(3)}{\bm{R}_{e}}\|$ where $\bm{R}_{e}=\bm{R}\bm{R}_{r}^{-1}$ and $\bm{R}_{r}=\bm{R}_{z}(\alpha_{d})\bm{R}_{y}(\beta_{d})\bm{R}_{z}(\gamma_{d})$. The results of the HIL experiments on fixed-wing UAV are presented in Figures 11 and 12.
Despite the errors introduced by the simplification in the derivation of setpoints and dynamic uncertainties, the control inputs can be effectively tracked by the fixed-wing UAV. This is because the proposed algorithm addresses both the curvature constraint and the nonholonomic constraint, which are the most critical kinematic constraints for the UAV. Consequently, both the distance error $d_{e}$ and the attitude error $\|\log_{\mathrm{SO}(3)}{\bm{R}_{e}}\|$ in Figure 12 are stabilized as soon as the motion planning algorithm is activated. This confirms the convergence of the UAV to the target positive limit set and its arrival at the target configuration under curvature constraint, as shown in Figure 11. Moreover, the results also demonstrate that the proposed algorithm can be implemented onboard in real-time, and can handle time-varying tasks where the target configuration switches during the flight, which is another advantage of the closed-form motion planning method.
VI-C Parameter Selection for Different Robots
In this subsection, we explain the effect of the curvature constraint in the construction of the CVF and provide a practical strategy for radii selection in the CVF on different robot platforms.
The theoretical foundation for this parameter selection strategy is established in Theorems 1 and 4, which provide a range of feasible values for the region radii, rather than designating specific values. This flexibility allows the region radii to be conveniently adjusted according to real-world scenarios, promising a wider applicability of the proposed method. For robots with better maneuverability and small turning radius (e.g., ground vehicles, unicycle-type robots), smaller and tighter region radii can be employed to create more compact CVFs with narrower transition regions. This enables faster convergence to the target positive limit set while maintaining desired curvature with a prescribed bound. In contrast, for robots with stricter dynamic constraints and larger turning radius (e.g., fixed-wing UAVs, large marine vessels), their CVFs are more sensitive to the change of control inputs. Therefore, larger and more disparate region radii are preferred to improve the robot’s tracking performance with smoother control inputs.
The experimental validation in this section demonstrates the above principle. In Section VI-A, the CVF for the Ackermann UGV with $\rho=$0.6\text{\,}\mathrm{m}$\$ is constructed with compact parameters $r_{1}=4\rho=$2.4\text{\,}\mathrm{m}$\$, $r_{2}=8\rho=$4.8\text{\,}\mathrm{m}$\$, $r_{3}=12\rho=$7.2\text{\,}\mathrm{m}$\$. It can be seen that the CVF evolves from source and sink to vortex in a narrower region with width $r_{3}-r_{1}=8\rho=$4.8\text{\,}\mathrm{m}$\$ as shown in Figure 13(a). In Section VI-B, the fixed-wing UAV with $\rho=$30\text{\,}\mathrm{m}$\$ requires more conservative parameters specified by $r_{1}=8\rho=$240\text{\,}\mathrm{m}$\$, $r_{2}=12\rho=$360\text{\,}\mathrm{m}$\$, $r_{3}=18\rho=$540\text{\,}\mathrm{m}$\$ to accommodate its sensitivity to control input variations under strict curvature constraints. The CVF with larger region radii and wider transition region is shown in Figure 13(b). The above parameter selection strategy ensures that the CVF-based motion planning is both theoretically sound and practically implementable across diverse robotic platforms.
(a) CVF for Ackermann UGV with minimum turning radius $\rho=$0.6\text{\,}\mathrm{m}$\$
(b) CVF for fixed-wing UAV with minimum turning radius $\rho=$30\text{\,}\mathrm{m}$\$
Figure 13: Comparison of the CVF for the different robots in hardware experiments, where the singular point $\bm{p}_{\delta}=\left[x_{\delta},y_{\delta}\right]^{\intercal}$ is presented by the red cross.
@@ Appendix A Proof of Lemma 1
as provided in [Stewart2011Calculus, Chapter 13.3], where the cross product in $\mathbb{R}^{2}$ is defined as $a,b]\times[c,d]\coloneq ad-bc$.
In the polar coordinates, the integral curve satisfies the following ordinary differential equation:
$\mathbf{F}=\left[\begin{matrix}F_{r}\\ F_{\varphi}\end{matrix}\right]=\left[\begin{matrix}\dot{r}\\ r\dot{\varphi}\end{matrix}\right].$ (32)
Moreover, the integral curve equation in Cartesian coordinates can be transformed from the polar coordinates into
$\bm{p}$ $=\left[\begin{matrix}r\cos{\varphi}\\ r\sin{\varphi}\end{matrix}\right]=R(\varphi)\left[\begin{matrix}r\\ 0\end{matrix}\right],$ (33)
$\dot{\bm{p}}$ $=\begin{bmatrix}F_{x}\\ F_{y}\end{bmatrix}=R(\varphi)\begin{bmatrix}F_{r}\\ F_{\varphi}\end{bmatrix}$,
where the 2-D rotation matrix $R(\varphi)$ is defined as
$R(\varphi)=\left[\begin{matrix}\cos{\varphi}&-\sin{\varphi}\\ \sin{\varphi}&\cos{\varphi}\end{matrix}\right].$ (34)
In light of the above equations, we further derive the second-order derivative of $\bm{p}$ as follows:
$\ddot{\bm{p}}$ $=\frac{\mathrm{d}}{\mathrm{d}t}(R(\varphi))\begin{bmatrix}F_{r}\\ F_{\varphi}\end{bmatrix}+R(\varphi)\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}F_{r}\\ F_{\varphi}\end{bmatrix}$ (35)
$=R(\varphi)\begin{bmatrix}0&-\dot{\varphi}\\ \dot{\varphi}&0\end{bmatrix}\begin{bmatrix}F_{r}\\ F_{\varphi}\end{bmatrix}+R(\varphi)\begin{bmatrix}\dot{F}_{r}\\ \dot{F}_{\varphi}\end{bmatrix}$
$=R(\varphi)\begin{bmatrix}0&-\frac{F_{\varphi}}{r}\\ \frac{F_{\varphi}}{r}&0\end{bmatrix}\begin{bmatrix}F_{r}\\ F_{\varphi}\end{bmatrix}$
$\quad+R(\varphi)\begin{bmatrix}\frac{\partial F_{r}}{\partial r}&\frac{\partial F_{r}}{r\partial\varphi}\\ \frac{\partial F_{\varphi}}{\partial r}&\frac{\partial F_{\varphi}}{r\partial\varphi}\end{bmatrix}\begin{bmatrix}\dot{r}\\ r\dot{\varphi}\end{bmatrix}$
$=R(\varphi)\begin{bmatrix}\frac{\partial F_{r}}{\partial r}&\frac{\partial F_{r}}{r\partial\varphi}-\frac{F_{\varphi}}{r}\\ \frac{\partial F_{\varphi}}{\partial r}+\frac{F_{\varphi}}{r}&\frac{\partial F_{\varphi}}{r\partial\varphi}\end{bmatrix}\mathbf{F}.$
By substituting 33 and 35 into (31), we can obtain (4).
@@ Appendix B Proof of Theorem 1
2) To prove that $\mathcal{O}_{2}$ is an almost stable limit cycle for integral curves with kinematics in Definition 5, we firstly show the existence of the limit cycle and then analyze its stability. Consider a closed annular region $\mathcal{A}_{inv}=\{\bm{p}\mid\|\bm{p}\|\in[\delta_{1},\ \delta_{2}],\ 0<\delta_{1}<r_{2},\ \delta_{2}>r_{2}\}$ containing $\mathcal{O}_{2}$ and denote the inner and outer boundaries of $\mathcal{A}_{inv}$ as $\mathcal{I}_{inv}$ and $\mathcal{O}_{inv}$, respectively. Since both boundaries are smooth, the tangent cone of $\mathcal{A}_{inv}$ in $\bm{p}\in\partial\mathcal{A}_{inv}$ is the tangent halfspace of $\partial\mathcal{A}_{inv}$ depicted by
@@
Given (10), it can be verified that $\mathbf{F}(\bm{p})\in\mathcal{C}_{\mathcal{A}_{inv}}(\bm{p})$, $\forall\bm{p}\in\partial\mathcal{A}_{inv}$, implying $\mathcal{A}_{inv}$ is positively invariant according to Nagumo’s theorem [nagumo1942lage]. Additionally, the positively invariant annular region $\mathcal{A}_{inv}$ contains at least one limit cycle since it contains no singular point and $\left<\mathbf{F}(\bm{p}),\bm{p}\right>\neq 0$, $\forall\bm{p}\in\partial\mathcal{A}_{inv}$ indicates the existence of forward orbit from the boundary [chicone2006ordinary, Theorem 1.179].
@@
According to (36a), there hold $\dot{r}_{d}>0$ for $r_{2}<r_{d}<0$ and $\dot{r}_{d}<0$ for $r_{d}>0$. Subsequently, we choose a positive definite Lyapunov function candidate:
@@
3) Since the VF $\mathbf{F}$ is of class $C^{1}$, we can conclude that the curvature is continuous over $\mathcal{D}$ according to Lemma 1. To show the boundedness of the curvature, we first discuss the trivial case where $\bm{p}\in\mathcal{A}_{1}\setminus\bm{0}$. In this case, there holds that $F_{r}=1$ and $F_{\varphi}=0$ and $\bm{K}_{F}=\bm{0}$. Substituting these values into (4), we have $\kappa_{\mathbf{F}}(\bm{p})=0$, $\forall\bm{p}\in\mathcal{A}_{1}\setminus\bm{0}$. Similarly we can obtain the same result that $\kappa_{\mathbf{F}}(\bm{p})=0$, $\forall\bm{p}\in\mathcal{A}_{4}$, where $F_{r}=-1$ and $F_{\varphi}=0$.
Next, consider $\bm{p}\in\mathcal{A}_{2}$, where $r_{1}\leq r<r_{2}$, $F_{r}=\lambda(s_{2})$, $F_{\varphi}=\tilde{\lambda}(s_{2})$, $\lambda(s_{2})=2s_{2}^{3}-3s_{2}^{2}+1$ and $s_{2}=(r-r_{1})(r_{2}-r_{1})\in[0,1)$. Substituting the above equations into (4) yields
@@
$=\lvert\frac{k_{1}(s_{2})}{r}+\frac{k_{2}(s_{2})}{r_{2}-r_{1}}\rvert\stackrel{{\scriptstyle r\geq r_{1}}}{{\leq}}\lvert\frac{k_{1}(s_{2})}{r_{1}}\rvert+\lvert\frac{k_{2}(s_{2})}{r_{2}-r_{1}}\rvert$
@@
$=\lvert\frac{k_{3}(s_{3})}{r}+\frac{k_{4}(s_{3})}{r_{3}-r_{2}}\rvert\stackrel{{\scriptstyle r\geq r_{2}}}{{\leq}}\lvert\frac{k_{3}(s_{3})}{r_{2}}\rvert+\lvert\frac{k_{4}(s_{3})}{r_{3}-r_{2}}\rvert$
@@ Appendix C Proof of Theorem 2
In this proof, we show that the proposed CVF has a bounded and continuous curvature and the target position $\bm{p}_{d}$ belongs to the positive limit set of integral curves and $\mathbf{T}(\bm{p}_{d})/\|\mathbf{T}(\bm{p}_{d})\|=[\cos{\theta_{d}},\sin{\theta_{d}}]^{\intercal}$.
1) Upper bounded continuous curvature.
According to [chicone2006ordinary, Chapter 1.5], the normalization operation reparameterizes the time of the integral curves of $\mathbf{T}$ from $\mathbf{F}$. In other words, the shape of integral curves of $\mathbf{T}$ with $\bm{\zeta}_{0}=\bm{p}$ remains the same as those of $\mathbf{F}$ with $\bm{\zeta}_{0}=\bm{p}+\bm{p}_{\delta}$. Therefore, the integral curves from $\mathbf{F}$ to $\mathbf{T}$ are translated by $\bm{p}_{\delta}$. In addition, the translation is a positive isometry, which preserves the curvature of the integral curves [banchoff2016Chap3.4]. Therefore, the integral curves of $\mathbf{T}$ still have continuous curvature bounded by $\bar{\kappa}$ as shown in Theorem 1.
2) Positive limit set of integral curves of $\mathbf{T}$.
@@
3) $\mathbf{T}(\bm{p}_{d})$ aligning with $\theta_{d}$.
Given (15), we transform the components of $\mathbf{F}$ from $(r,\varphi)$ coordinates into Cartesian coordinates by rotation matrix $R(\varphi)$ and there holds that
@@
$=\lim_{\begin{subarray}{c}r\to r_{2}\\ \varphi\to\theta_{d}-\frac{\pi}{2}\end{subarray}}R(\varphi)\left[\begin{matrix}F_{r}(r)\\ F_{\varphi}(r)\end{matrix}\right]/\|R(\varphi)\left[\begin{matrix}F_{r}(r)\\ F_{\varphi}(r)\end{matrix}\right]\|$
@@ Appendix D Kinematics of Reference Orientation
$\varphi_{\delta}$ $=\arctan{\frac{\left<\bm{p}-\bm{p}_{\delta},\bm{e}_{y}\right>}{\left<\bm{p}-\bm{p}_{\delta},\bm{e}_{x}\right>}},$
@@
$\left[\begin{matrix}T_{x}\\ T_{y}\end{matrix}\right]=R(\varphi_{\delta})\left[\begin{matrix}T_{r_{\delta}}\\ T_{\varphi_{\delta}}\end{matrix}\right],$
@@
$\left[\begin{matrix}-T_{y}&T_{x}\end{matrix}\right]=\left[\begin{matrix}T_{r_{\delta}}&T_{\varphi_{\delta}}\end{matrix}\right]R{\left(-(\varphi_{\delta}+\frac{\pi}{2})\right)}.$
@@
$\dot{\mathbf{T}}=\left[\begin{matrix}\dot{T}_{x}\\ \dot{T}_{y}\end{matrix}\right]=\frac{\mathrm{d}}{\mathrm{d}t}\left(R{(\varphi_{\delta})}\left[\begin{matrix}T_{r_{\delta}}\\ T_{\varphi_{\delta}}\end{matrix}\right]\right)$
$=\frac{\mathrm{d}R{(\varphi_{\delta})}}{\mathrm{d}t}\left[\begin{matrix}T_{r_{\delta}}\\ T_{\varphi_{\delta}}\end{matrix}\right]+R{(\varphi_{\delta})}\frac{\mathrm{d}}{\mathrm{d}t}\left[\begin{matrix}T_{r_{\delta}}\\ T_{\varphi_{\delta}}\end{matrix}\right].$
@@
where $\Delta\theta=\theta-\theta_{\nabla}$, $\theta_{\nabla}=\arctan{\frac{\left<\nabla\theta_{r},\bm{e}_{y}\right>}{\left<\nabla\theta_{r},\bm{e}_{x}\right>}}$ and $A(r_{\delta})=\|\nabla\theta_{r}\|$ denote the direction and magnitude of the gradient $\nabla\theta_{r}$, respectively. Moreover, given the partial derivative of $\theta_{r}$ w.r.t. $r_{\delta}$:
@@
where $s_{i}=(r-r_{i-1})/(r_{i}-r_{i-1})$, $\lambda_{i}=2s^{3}_{i}-3s^{2}_{i}+1$ for $i=2,3$ and $\mathcal{A}_{i}$ are defined in (9), the direction and magnitude of $\nabla\theta_{r}$ can be computed by $\theta_{\nabla}=\varphi_{\delta}+\mathrm{arccot}(r_{\delta}\theta_{r,r_{\delta}})$ and $A(r_{\delta})=\sqrt{1/r_{\delta}^{2}+\theta_{r,r_{\delta}}^{2}}$, respectively.